### 08 Mar 2024 v1

# RadarDistill: Boosting Radar-based Object Detection Performance via Knowledge Distillation from LiDAR Features

Abstract
The inherent noisy and sparse characteristics of radar data pose challenges in finding effective representations for 3D object detection. In this paper, we propose RadarDistill, a novel knowledge distillation (KD) method, which can improve the representation of radar data by leveraging LiDAR data. RadarDistill successfully transfers desirable characteristics of LiDAR features into radar features using three key components: Cross-Modality Alignment (CMA), Activation-based Feature Distillation (AFD), and Proposal-based Feature Distillation (PFD). CMA enhances the density of radar features through multiple layers of dilation operations, effectively addressing the challenges of inefficient knowledge transfer from LiDAR to radar. AFD is designed to transfer knowledge from significant areas of the LiDAR features, specifically those regions where activation intensity exceeds a predetermined threshold. PFD guides the radar network to mimic LiDAR network features in the object proposals for accurately detected results while moderating features for misdetected proposals like false positives. Our comparative analyses conducted on the nuScenes datasets demonstrate that RadarDistill achieves state-of-the-art (SOTA) performance for radar-only object detection task, recording 20.5% in mAP and 43.7% in NDS. Also, RadarDistill significantly improves the performance of the camera-radar fusion model.
Figure 1: Illustration of RadarDistill. Through knowledge transfer from LiDAR features to radar features, our RadarDistill converts sparse and noisy radar features into dense and semantically rich LiDAR-like features for better object detection.
1 Introduction
While 3D perception based on camera and LiDAR sensors has been widely studied, radar sensors are now gaining attention due to their affordability and reliability in adverse weather conditions. Radar sensors can locate objects in a Bird’s Eye View (BEV) and also measure their radial velocity using Doppler frequency analysis. However, when compared to LiDAR or camera sensors, radar’s major limitations are its lower spatial resolution and a higher likelihood of false positives due to multi-path reflections. For decades, traditional object detection and tracking methods, based on manually crafted models, have been developed to overcome these limitations by many radar manufacturers. While deep neural networks (DNNs) have considerably improved 3D perception in camera and LiDAR sensors, similar advancements have not been mirrored in radar sensor-specific architectures. There are only a handful of studies that have applied deep neural networks to radar data. For instance, KPConvPillars [28] and Radar-PointGNN [26] leveraged KPConv [27] and graph neural networks, respectively, to detect objects using radar point clouds. However, these methods have not yet achieved the level of significant improvements realized with camera or LiDAR data. Recently, it was shown that radar can be effectively fused with camera or LiDAR data to enhance the robustness of 3D object detection [14, 15, 42, 21, 31, 13].
This paper focuses on improving the performance of radar-based 3D object detection using deep neural networks. We note that the limited performance of radar is largely due to the challenges in finding effective representations, given the sparse and noisy nature of radar measurements. Inspired by the remarkable success of deep models that encode LiDAR point clouds, our goal is to transfer the knowledge extracted from a LiDAR-based model to enhance a radar-based model. Recently, knowledge distillation (KD) techniques have shown success in transferring knowledge from one sensor modality to another, thereby refining the representation quality of the target model. To date, various KD methods have been introduced in the literature [8, 4, 11, 3, 41, 16, 32, 40]. Cross-modality knowledge distillation can be broadly divided into two approaches based on the modality of the student model used. When employing a camera as the student model, both depth and shape information are transferred from the teacher model to the student model. BEVDistill [3] transforms both LiDAR and cameras features into a BEV format, enabling the transfer of spatial knowledge from LiDAR features to camera features. DistillBEV [32] utilizes the prediction results from either LiDAR or LiDAR-camera fusion models to distinguish between foreground and background, guiding the student model to focus on KD in essential areas. S2M2-SSD [40] determines key areas based on the student model’s predictions and transfers information obtained from a LiDAR-camera fusion model in the key areas. Apart from these approaches, UniDistill [41] employs a universal cross-modality framework that enables knowledge transfer among diverse modalities. This framework is adaptable to different modality pairings, including camera-to-LiDAR, LiDAR-to-camera, and (camera+LiDAR)-to-camera settings.
In this paper, we propose RadarDistill, a novel KD framework designed to enhance the representation of radar data, leveraging LiDAR data. Our study shows that by employing a radar encoding network as a student network and a LiDAR encoding network as a teacher network, our KD framework effectively produces radar features akin to the dense and semantically rich features derived from LiDAR data for better object detection. Although both LiDAR data and its encoding network are used to enhance radar features during the training phase, they are not required in the inference phase.
Our proposed RadarDistill is designed based on three main ideas: 1) Cross-Modality Alignment (CMA), 2) Activation-based Feature Distillation (AFD), and 3) Proposal-based Feature Distillation (PFD). Our study indicates that transferring knowledge from LiDAR to radar features is difficult due to the inherent sparsity of radar data, which complicates the alignment with the more densely distributed LiDAR features. To address this problem, CMA boosts the student network’s capacity and simultaneously increases the ratio of activated radar features by implementing multiple layers of dilation operations.
The proposed AFD and PFD minimize the distribution gap between the intermediate features produced by the radar and LiDAR encoding networks. Initially, AFD conducts Activation-aware Feature Matching on low-level features. Specifically, it divides both radar and LiDAR features into active and inactive regions according to activation intensity of each features and constructs KD losses for each region separately. By assigning greater importance to the KD loss linked with active regions, AFD enables the network to concentrate on transferring knowledge for significant features.
Next, PFD implements Proposal-level Selective Feature Matching, aimed at narrowing the differences from features associated with the proposals generated by the radar detection head. PFD directs the radar network to generate object features that are similar in shape to the high-level LiDAR features for accurately detected proposals. Conversely, for misdetected proposals such as false positives, the model is guided to suppress falsely activated features, reflecting low activation of the LiDAR features.
[…]
The key contributions of our work are as follows:
  • Our study is the first to demonstrate that radar object detection can be substantially improved using LiDAR data during the training process. Our qualitative results in Fig. 1 highlight that the radar features acquired through RadarDistill successfully mimic those of LiDAR, leading to enhanced object detection and localization.
  • Our findings reveal that the CMA is a crucial element of RadarDistill. In its absence, we observed a significant drop in performance enhancement. According to our ablation study, CMA plays a pivotal role in resolving inefficient knowledge transfer caused by the different densities of radar and LiDAR point clouds.
  • We propose two novel KD methods, AFD and PFD. These methods bridge the discrepancy between radar and LiDAR features, operating at two separate feature levels and utilizing KD losses specifically designed for each level.
  • RadarDistill achieves the state-of-the-art performance in the radar-only object detector category on the nuScenes benchmark. It also achieves a significant performance boost for camera-radar fusion scenarios.
2 Related Works
2.1 Radar-based 3D Object Detection
Radar-only 3D object detection models have employed backbone models adopted from various LiDAR-based detectors to suit the specific needs of radar data. Radar-PointGNN [26] utilized GNNs [24] to effectively extract features from sparse radar point clouds through a graph representation. KPConvPillars [28] introduced a hybrid architecture that combined grid-based and point-based approaches for radar-only 3D object detection. Recent studies have concentrated on detection techniques that combine radar with LiDAR or cameras [21, 14, 15, 42, 34, 31]. These methods aimed to complement the limitations of each sensor by using the data obtained from radars. RadarNet [34] employed a voxel-based early fusion and an attention-based late fusion approach to integrate radar and LiDAR data. RCM-Fusion [13] fused radar and camera data at both the feature level and instance level to fully utilize the potential of radar information. CRN [15] leveraged radar data to transform camera image features into a BEV view and then combined these transformed camera features with the radar BEV features using multi-modal deformable attention.
2.2 Knowledge Distillation
Knowledge distillation, as a strategy for model compression, transfers the information from a larger teacher model with greater capacity to a smaller student model. The student model mimics the teacher’s intermediate features [22], prediction logit [10], or activation boundary [9] to acquire knowledge from the teacher model. KD approaches have recently been extended from image classification to object detection [2, 17, 29, 7, 37, 35]. Defeat [7] utilized KD by decoupling features to foreground and background with ground truth. FGD [37] employed spatial and channel attention for 2D object detection, guiding the student model to focus on critical pixels and channels.
In 3D object detection, KD was applied to reduce model complexity in 2D detection [36, 19, 39], or to enhance detection performance through cross-modality knowledge distillation [8, 4, 11, 18, 40, 3, 41, 16, 32]. SparseKD [36] focused on KD for primary regions identified by the teacher model’s predictions, and the student model was initialized with the weights of the teacher to enhance feature extraction. MonoDistill [4] projected LiDAR point clouds onto the image plane and used a 2D CNN to develop an ‘image-version’ LiDAR-based model, effectively bridging the gap in feature representations between LiDAR and camera. BEVDistill [3] projected LiDAR and camera features into the BEV space, effectively transferring 3D spatial knowledge through dense feature distillation and sparse instance distillation. UniDistill [41] proposed a universal framework that was suitable for various teacher-student pairs. UniDistill projected both teacher and student features into BEVs and applied KD on the object features within each ground truth box.
As compared to the existing studies on cross-modality KD methods, our study focuses on developing KD specifically tailored for radar object detection, considering the sparse and noisy nature of radar data.
Figure 2: Overall architecture of RadarDistill. The input point clouds from each modality are initially processed through Pillar Encoding and Sparse 2D CNN to extract low-level features. CMA processes the low-level features of the radar to make them denser and align them with the features of LiDAR. AFD distinguishes between active and inactive regions in both radar and LiDAR features and minimizes their KD losses using different importance factors. Subsequently, proposal-level features are identified using the detection head, and PFD reduces the distribution gap between proposal features obtained from the radar and LiDAR networks. LiDAR data is only needed during the training phase to enhance radar features and not during inference.
3 Method
In this section, we present the details of the proposed KD framework, RadarDistll. Fig. 2 illustrates the overall architecture of RadarDistill. In Section 3.1, we provide a brief overview of the baseline model employed as both teacher and student models. In Section 3.2, we describe the details of Cross-Modality Alignment, a module for densifying radar features. In Section 3.3, we present the AFD, a novel Activation-aware Feature Matching approach. Finally Section 3.4 outlines the PFD, a Proposal-level Selective Feature Matching approach.
3.1 Preliminary
We used PillarNet [23] as our baseline model for both LiDAR and radar detectors. PillarNet initially generates regular pillars from the point clouds, subsequently producing sparse 2D pillar features denoted as $F^{2D}_{ldr}$ for LiDAR and $F^{2D}_{rdr}$ for radar through the pillar feature network. Then, we define the low-level BEV features denoted by $F^{l}_{ldr}$ and $F^{l}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$ obtained by encoding 2D pillar features with a 2D sparse convolution-based encoder (SparseEnc), as follows
$F^{l}_{mod}=SparseEnc(F^{2D}_{mod}),$ (1)
where low-level BEV features are sparse features derived from processing through SPConv [6] and $mod$ represents the detector modality, either ldr or rdr. We describe the high-level BEV features denoted by $F^{h}_{ldr}$ and $F^{h}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$, formed by encoding low-level BEV features $F^{l}_{ldr}$ and $F^{l}_{rdr}$ using a 2D dense convolution-based encoder (DenseEnc) as
$F^{h}_{mod}=DenseEnc(F^{l}_{mod}).$ (2)
The high-level BEV features are processed through a centerpoint[38]-based head (CenterHead) to generate the final prediction heatmaps for classification, regression, and IoU scoring, represented as $H^{cls}$, $H^{reg}$ and $H^{IoU}$. The final predictions are obtained from
$H^{cls}_{mod},H^{reg}_{mod},H^{IoU}_{mod}=CenterHead(F^{h}_{mod}).$ (3)
3.2 Cross-Modality Alignment
Our goal is to reduce the disparity between radar and LiDAR due to the differences in point cloud density via CMA. PillarNet [23] employs SPConv [6] to generate low-level features from only non-empty pillars. Notably, when comparing the number of non-empty pillars, radar’s non-empty pillars ($F^{l}_{rdr}$) comprise only $11\%$ of the total count of LiDAR’s non-empty pillars ($F^{l}_{ldr}$). This significant disparity in the quantity of non-empty pillars necessitates an alignment of the two modalities, especially since the information from LiDAR’s non-empty pillars cannot be directly transferred to the corresponding empty pillars in radar data.
Fig. 3 presents the structure of the CMA. Within the component of CMA, the down block applies deformable convolution [5] to selectively extract essential features during downsampling, and then it aggregates these features using the ConvNeXt V2 blocks [33]. In the up block, dilation is applied using 2D transposed convolution to densify the features in surrounding areas. The aggregation module that includes concatenation followed by a 1$\times$1 convolution layer combines different features and operates similarly to a skip connection.
We downsample the 1/8 resolution features $F^{l}_{rdr}$ to produce $F^{l^{\prime(1)}}_{rdr}$. These $F^{l^{\prime(1)}}_{rdr}$ features are upsampled and combined with $F^{l}_{rdr}$ in the aggregation module, forming the aligned BEV features $F^{a^{(1)}}$. Subsequently, we reduce the resolution of $F^{l^{\prime(1)}}_{rdr}$ to 1/32 and then increase it to 1/16. These upsampled features are integrated with the downsampled $F^{a^{(1)}}$, to produce $F^{l^{\prime(2)}}_{rdr}$. Finally, we upsample the $F^{l^{\prime(2)}}_{rdr}$ and fuse it with $F^{a^{(1)}}$ in the aggregation module, which generates the second aligned BEV features $F^{a^{(2)}}$. In the CMA process, we acquire two sets of aligned features at 1/8 resolution. We apply multiple AFD to these aligned features to enhance their similarity to LiDAR’s low-level features. Consequently, as observed in Fig. 1, the radar’s low-level features become denser after being processed through the CMA. A comprehensive description of the CMA’s component is available in the Supplementary Materials.
3.3 Activation-based Feature Distillation
AFD strategically aligns the low-level features of radar and LiDAR through Activation-aware Feature Matching. This process matches the activation patterns of radar features to those of LiDAR, thus bridging the gap in their distribution of active features. Additionally, AFD is applied iteratively to each pair of the $n$-th aligned features, $F^{a^{(n)}}$ and $F^{l}_{ldr}$, ensuring a more precise and effective alignment between the radar and LiDAR features. First, for both LiDAR and radar features, we define “active regions” as areas where each radar and LiDAR feature intensity values, summed in the channel direction, are greater than 0. Conversely, areas where these sums are less than 0 are classified as “inactive regions”. We define active mask $M^{mod}_{i,j}$ to present these regions and obtain the active masks $M^{a^{(n)}}_{i,j}$and $M^{ldr}_{i,j}$ for $F^{a^{(n)}}$and $F^{l}_{ldr}$ correspondingly. These masks are obtained from
$F^{mod}_{i,j}=\sum^{C}_{c=1}F^{mod}_{c,i,j},$ (4)
$M^{mod}_{i,j}=\begin{cases}1,&if\,(F^{mod}_{i,j}>0),\\ 0,&otherwise,\end{cases}$ (5)
where $F\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$ represents BEV features, the indices i,j,c correspond to height, width, and channel, and $mod$ indicates the modality of features, comprising either ldr or a${}^{(n)}$. The distillation region is adaptively separated into two categories: the active regions (AR) where both radar and LiDAR are active and the inactive regions (IR) where radar is active while LiDAR is inactive. There is an imbalance between the AR and IR due to the sparse and noisy characteristics of radar. This can disturb training by focusing on imitating the dominant IR. Therefore, we apply a relative adaptive weight based on the number of pixels in each region. AR is assigned a weight of 1 as it is a smaller yet crucial region while IR is given a relative adaptive weight denoted by $\rho$ as determined by the ratio of pixel counts between AR and IR. We define a separate mask $M^{sep^{(n)}}$ for distinguishing the regions, specifically between $F^{a^{(n)}}$and $F^{l}_{ldr}$ as
$AR^{(n)}$ $=\,(M^{a^{(n)}}=1)\,\&\,(M^{ldr}=1),$ (6)
$IR^{(n)}$ $=\,(M^{a^{(n)}}=1)\,\&\,(M^{ldr}=0),$ (7)
$M^{sep^{(n)}}_{i,j}$ =$=\begin{cases}\alpha,&if\,(i,j)\in AR^{(n)},\\ \rho^{(n)}\times\beta,&if\,(i,j)\in IR^{(n)},\\ 0,&\,otherwise,\end{cases}$ (8)
$\rho^{(n)}$ $=\frac{N_{AR^{(n)}}}{N_{IR^{(n)}}},$ (9)
where $\rho^{(n)}$ represents the relative importance of the $IR^{(n)}$ in comparison to $AR^{(n)}$, $\alpha$ and $\beta$ are the balancing parameters between the separated regions, and $N$ is the number of pixels. In this way, the distillation region and the $\rho^{(n)}$ value are adaptively adjusted for each training, and distillation is performed with different weights for AR and IR.
Figure 3: Detailed structure of the proposed CMA module.
After decomposing the distillation region with AR and IR, the $n$-th aligned features $F^{a^{(n)}}_{rdr}$ are trained to imitate $F^{l}_{ldr}$. The distillation loss $L^{(n)}_{low}$ for the $n$-th activation features is given by
$L^{(n)}_{low}$ $=\sum^{C}_{c=1}\sum^{H}_{i=1}\sum^{W}_{j=1}M^{sep^{(n)}}_{i,j}(F^{ldr}_{c,i,j}-F^{a^{(n)}}_{c,i,j})^{2},$ (10)
where $F^{ldr}$ denotes the LiDAR’s low-level BEV features $F^{l}_{ldr}$.
The active mask loss is applied between the predicted mask, obtained by applying a sigmoid function to $F^{a(n)}$, and $M^{ldr}$ to ensure the radar’s active mask closely mimics LiDAR. The active mask loss $L^{(n)}_{mask}$ is calculated as
$L^{(n)}_{mask}$ $=\sum^{H}_{i=1}\sum^{W}_{j=1}|M^{ldr}_{i,j}-\sigma(F^{a^{(n)}}_{i,j})|,$ (11)
where $\sigma(\cdot)$ denotes a sigmoid function.
We apply AFD twice to the aligned BEV features to more accurately imitate $F^{l}_{ldr}$, and use the average of losses calculated from these multiple AFD as the final AFD loss. The final AFD loss function is given by
$L_{AFD}$ $=\frac{1}{2}\sum^{2}_{n=1}(L^{(n)}_{low}+L^{(n)}_{mask}).$ (12)
3.4 Proposal-based Feature Distillation
PFD employs Proposal-level Feature Matching to reduce the discrepancies between the high-level features. It utilizes predictions generated from the radar, guiding the radar network to create object features that closely mimic the high-level features of LiDAR, thereby achieving a more accurate alignment in properly identified proposals. Additionally, for erroneously identified proposals, such as false positives, PFD directs the model to suppress falsely activated features, thereby mirroring the lower activation levels typically seen in LiDAR features for such instances.
We employ the radar’s predicted heatmap to adaptively define the crucial region according to radar. Additionally, there is a significant scale disparity between the high-level BEV features generated by radar and LiDAR. PFD aligns the scales between the two modalities by making the student mimic the normalized values that are applied in the channel direction for both radar and LiDAR. DenseEnc perform input $F^{l}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$ by first downsampling to 1/16 of its size and then upsampling to 1/8 to generate the initial high-level BEV feature $F^{h^{(1)}}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$. Subsequently, $F^{h^{(1)}}_{rdr}$ and CMA’s output $F^{a^{(2)}}_{rdr}$ are concatenated and then fed through a convolution layer resulting in the second high-level BEV feature $F^{h^{(2)}}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$. Similar to AFD, PFD is applied twice to each of the multiple feature sets $F^{h^{(m)}}_{rdr}$.
The first step involves thresholding the predicted radar heatmap $H^{cls}_{rdr}$ and the GT heatmap $H^{cls}_{GT}$ to identify the regions of true positives, false positives, and false negatives. The regions are identified using the following operations
$TP$ $=\,(H^{cls}_{GT}>\sigma)\,\&\,(H^{cls}_{rdr}>\sigma),$ (13)
$FP$ $=\,(H^{cls}_{GT}<\sigma)\,\&\,(H^{cls}_{rdr}>\sigma),$ (14)
$FN$ $=\,(H^{cls}_{GT}>\sigma)\,\&\,(H^{cls}_{rdr}<\sigma),$ (15)
where $H^{cls}_{rdr}$ represents the predicted radar classification heatmap, $H^{cls}_{GT}$ is the GT classification heatmap, and $\sigma$ is the threshold parameter set at 0.1.
TP and FN are considered important regions according to the foreground, while FP indicates regions where radar is mostly wrong. Consequently, TP and FN are categorized into one group, FP into another, leading to the generation of a scaling mask $M^{scale}$ for each defined area. The scaling mask is expressed as
$M^{scale}_{i,j}$ =$=\begin{cases}\frac{\lambda_{1}}{N_{TP}+N_{FN}},&if\,(i,j)\in(TP\cup FN),\\ \frac{\lambda_{2}}{N_{FP}},&if\,(i,j)\in FP,\\ 0,&\,otherwise,\end{cases}$ (16)
where $N$ denotes the number of pixels, while $\lambda_{1}$ and $\lambda_{2}$ are balancing parameters for the respective regions. The alignment of feature scales between radar and LiDAR involves applying the channel-wise softmax to the high-level BEV features $F^{h}_{ldr}$ and $F^{h^{(m)}}_{rdr}$, which results in $S^{ldr}$ and $S^{h^{(m)}}$, i.e.,
$S^{mod}_{c,:,:}=\frac{{\text{e}xp}(\frac{F^{mod}_{c,:,:}}{\tau})}{\sum^{C}_{c^{*}=1}{\text{e}xp}(\frac{F^{mod}_{c^{*},:,:}}{\tau})},$ (17)
where $mod\in\{ldr,h^{(m)}\}$, $\tau$ is a hyper-parameter that controls the smoothing of the distribution and is used as 0.5. Similar to AFD, PFD is employed twice to train $S^{h^{(m)}}$ to imitate $S^{ldr}$ and is computed as
$L^{(m)}_{high}=\sum^{C}_{c=1}\sum^{H}_{i=1}\sum^{W}_{j=1}M^{scale}_{i,j}|S^{ldr}_{c,i,j}-S^{h^{(m)}}_{c,i,j}|.$ (18)
The final PFD loss value is derived from the average of the PFD loss values computed from multiple PFD. The final PFD loss is calculated from
$L_{PFD}=\frac{1}{2}\sum^{2}_{m=1}(L^{(m)}_{high}).$ (19)
3.5 Overall Loss Function
We freeze the teacher model that has been pre-trained on LiDAR point clouds and then train the student model end-to-end using radar point clouds. Finally, we compute the total loss as the weighted sum of the detection loss and the distillation losses as
$L_{total}=L_{det}+\gamma L_{AFD}+\delta L_{PFD},$ (20)
where $\gamma$ and $\delta$ are hyperparameters used to balance the scale of different distillation losses.
4 Experiments
4.1 Experimental Setup
Datasets and metrics. We conduct the experiments on the nuScenes [1] dataset. This dataset contains 700, 150, 150 driving scenes for training, validation, and testing each captured with images from six cameras, points from five radars and one LiDAR. Our evaluation utilizes the official metrics from nuScenes, which include mean average precision (mAP) and the nuScenes detection score (NDS).
[…]
Radar-PointGNN [26] R - 0.5 3.4 1.024 0.859 0.897 1.020 0.931
KPConvPillars [28] R - 4.9 13.9 0.823 0.428 0.607 2.081 1.000
PillarNet* [23] (Our baseline) R - 8.6 34.7 0.532 0.283 0.615 0.438 0.092
  RadarDistill R 20.5 43.7 0.461 0.263 0.525 0.336 0.072
Table 1: Performance comparison on nuScenes test set. The best performed metrics in the radar-only model are bolded. ”*” denotes models we have reproduced without applying test time augmentation. Our model achieves the state-of-the-art performance in all metrics.
Method Input KD Car Truck bus Trailer C.V Ped. Motor. Bicycle T.C Barrier
CenterFusion [21] C, R - 50.9 25.8 23.4 23.5 7.7 37.0 31.4 20.1 57.5 48.4
FCOS3D [30] C - 52.4 27.0 27.7 25.5 11.7 39.7 34.5 29.8 55.7 53.8
MonoDIS [25] C - 47.8 22.0 18.8 17.6 7.4 37.0 29.0 24.5 48.7 51.1
CenterNet [43] C - 53.6 27.0 24.8 25.1 8.6 37.5 29.1 20.7 58.3 53.3
PillarNet* [23] (Our baseline) R - 41.8 11.6 8.4 6.5 0.0 7.4 1.0 0.0 1.1 8.1
  RadarDistill R 54.0 15.3 11.3 29.5 5.5 9.2 15.3 0.9 21.7 42.3
Table 2: Per-class comparison on nuScenes test set. The best performed metrics are bolded. ’C.V’, ’Ped.’, ’Motor.’, and ’T.C.’ represent construction vehicle, pedestrian, motorcycle, and traffic cone, respectively.
Implementation details. As a baseline model, we utilized PillarNet-18, which is PillarNet [23] with a ResNet-18 backbone. We used the Adam optimizer with a learning rate of 0.001, applying a one-cycle learning rate policy. We set weight decay at 0.01, and adjusting momentum between 0.85 and 0.95. We trained our baseline models on 4 NVIDIA RTX 3090 GPUs for 20 epochs, employing a total batch size of 16 with the Class-Balanced Grouping and Sampling (CBGS) strategy [44]. We applied random scene flipping along the axes, random rotation, scaling, translation and ground-truth box sampling as the data augmentation. We set our detection range on the $X$ and $Y$ axes to $[-54.0m,54.0m]$ and on the Z-axis to $[-3m,5m]$ , utilizing pillars of $(0.075m,0.075m)$. Our proposed model was trained for $40$ epochs and all other training procedures were applied identically to the baseline. We adopted the teacher’s weights to initialize the student model, following the inheriting strategy [12]. In Eq. (9), we set the hyper-parameters $\alpha$ to 3e-4 and $\beta$ to 5e-5, while in Eq. (16), the hyper-parameters $\lambda_{1}$ and $\lambda_{2}$ were set to $5$ and $1$, respectively. Additionally, in Eq. (20), we set the parameter $\gamma$ and $\delta$ to $5$ and $25$.
4.2 Comparison with the State-of-the-Arts
We evaluated the performance of RadarDistill on nuScenes test set. In Table 1, our RadarDistill achieved top-ranking performance on the nuScenes radar-only benchmark, significantly outperforming other methods. Specifically, our model achieved a $20.5\%$ mAP and $43.7\%$ NDS, surpassing the previous state-of-the-art (SOTA) method, KPConvPillars [28], by a substantial margin of $+15.6\%$ in mAP and $+29.8\%$ in NDS.
As shown in Table 2, we compared the performance of our RadarDistill for each class with other camera-only and camera-radar fusion models. Our model outperformed the baseline model (PillarNet [23]), especially for the barrier and trailer classes, with gains of at least $+23\%$. Furthermore, our model demonstrated the best performance for the car and trailer classes even when compared with other camera-only models and camera-radar fusion model approaches. However, the pedestrian and bicycle classes showed only slight performance improvements, largely due to the inherent limitations of radar sensors, as these objects are smaller and have few radar points in their ground truth (GT) boxes.
4.3 Ablation Studies
We conducted several ablation studies using the nuScenes validation set. We trained our model on 1/7 of training set and used the complete validation set for evaluation to reduce the time spent on experiments.
Component Analysis. We present an ablation study to evaluate the impact of each component of our proposed model, as shown in Table 3. In #2, when CMA is applied to the baseline, there’s an overall enhancement in performance; yet, the improvements are relatively low. Comparing #2 with #3 and #4, where CMA is combined with AFD and PFD respectively, we see a significant performance gain. For #3, there is an increase in mAP by $+4.5\%$ and NDS by $+4.4\%$. In the case of #4, mAP improves by $+1.8\%$ and NDS by $+3.0\%$. Additionally, when all components are used together as in #6, we can see that it achieves the best performance across all metrics. Lastly, to understand the influence of CMA in the distillation process, #5 was tested without it. The results show a significant difference in performance, with an increase of $+4.2\%$ in mAP and $+4.0\%$ in NDS when CMA is included, demonstrating that the CMA structure performs a critical role in our model.
Comparative Analysis of Distillation Regions. We conducted experiments to analyze the importance of distillation regions during the knowledge distillation (KD) process.
Table 4 presents an ablation study conducted to evaluate the impact of selecting different regions when applying low-level feature distillation. When distillation is performed on the entire area without separate regions, we observe a performance improvement of $+4.2\%$ in Car AP, $+2.2\%$ in mAP, and $+2.9\%$ in NDS. Applying distillation to the gaussian heatmap areas centered around the GT boxes led to an increase in Car AP and mAP compared to the total, but a decrease in NDS. Dividing the distillation between foreground and background regions, compared to the total, resulted in an increase in both Car AP and mAP, but a slight decrease in NDS. Conversely, dividing the feature distillation by our proposed activation regions yielded the best performance across all metrics, with improvements of $+6.8\%$ in Car AP, $+2.3\%$ in mAP, and $+1.5\%$ in NDS compared to total.
Next, Table 5 presents an experiment comparing the performance between previous cross-modality KD methods and PFD’s approach in applying high-level feature distillation. Here, the first row is based on the application of AFD according to our proposed method. When distillation is performed on the entire area, we observe a slight decrease in Car AP and mAP by $-0.6\%$ and $-0.1\%$ respectively, while NDS shows a marginal increase of $0.1\%$, indicating a minimal change in performance. Even when comparing the gaussian heatmap areas centered around the GT boxes, only Car AP shows a slight increase $0.1\%$ with the rest being identical to the total. Furthermore, applying KD based on LiDAR prediction results also did not yield any notable performance improvements. In contrast, our method, which applies KD based on radar prediction boxes, showed an improvement of over $0.4\%$ across all metrics compared to the total, demonstrating the best performance.
Scale Alignment in PFD. We conducted experiments to examine the impact of aligning the scale between teacher and student in PFD process in Table 6. When applying PFD without normalization, we observe a substantial decrease of $-5.9\%$ in Car AP, $-2.4\%$ in mAP, and $-1.9\%$ in NDS compared to before applying PFD. Conversely, after implementing normalization and PFD, we observe an increase of $+0.3\%$ in Car AP, $+0.3\%$ in mAP, and $+1.0\%$ in NDS.
Apply radar-camera fusion model. Table 7 shows a comparison of the performance when using radar features enhanced by our method in the radar-camera fusion model. The enhanced radar and camera features were fused using the BEVFusion [20] method, with its detailed structure and training method provided in the Supplementary Materials. We observe a significant increase of $+1.8\%$ in Car AP, $+1.3\%$ in mAP, and $+1.1\%$ in NDS.
[…]
1 5.4 27.3
2 6.4 29.3
3 10.9 33.7
4 8.2 32.3
5 7.0 30.7
6 11.2 34.7
Table 3: Ablation study of components in Radardistill. “CMA”: Cross-Modality Alignment. “AFD”: Activation-based Feature Distillation. “PFD”: Proposal-based Feature Distillation.
Region AFD
[…]
W/O AFD 34.8 6.4 29.3
Total 39.0 8.6 32.2
GT 40.4 8.8 29.5
FG / BG 42.1 9.6 31.7
Activation 45.8 10.9 33.7
Table 4: Ablation study of distillation regions in AFD. “W/O AFD”: Applying only CMA without distillation. “FG/BG”: Foreground and background region.
Region PFD
[…]
W/O PFD 45.8 10.9 33.7
Total 45.2 10.8 33.8
GT 45.9 10.8 33.8
LiDAR prediction 45.4 10.8 34.0
Radar prediction 46.1 11.2 34.7
Table 5: Ablation study of distillation regions in PFD. “W/O PFD”: Applying only CMA and AFD. “LiDAR prediction” : Using only LiDAR prediction results. “Radar prediction” : Using only Radar prediction results
[…]
W/O PFD 45.8 10.9 33.7
W/O Norm 39.9 8.5 31.8
Norm 46.1 11.2 34.7
Table 6: Ablation study of scale alignment in PFD. “W/O PFD”: Applying only CMA and AFD. “W/O Norm”: Applying PFD without normalizing. “Norm”: Applying normalizing and PFD.
[…]
BEVFusion* [20] C,R 65.9 38.3 45.3
RadarDistill-CR C,R 67.7 39.6 46.4
Table 7: Ablation study of applying to a radar-camera fusion model. ”*” denotes models we have reproduced without applying test time augmentation. Our model achieves slight performance enhancements in all metrics.
5 Conclusion
In our study, we introduced a new radar-based 3D object detection method, RadarDistill, which leveraged LiDAR data via knowledge distillation for enhanced performance. Our approach successfully guided the radar encoding network to generate radar features that closely resemble the detailed and semantically rich features of LiDAR. CMA enhanced the student network’s expressive power beyond that of the teacher network, allowing the student network to learn the complex distribution of LiDAR features more effectively. Through Activation-aware Feature Matching and Proposal-level Selective Feature Matching, both AFD and PFD successfully reduced the discrepancy from LiDAR feature distributions, particularly in important regions within the intermediate features. Our experiments demonstrated that RadarDistill achieved the state-of-the-art performance in radar-only object detection and offered significant performance gains in camera-radar fusion scenarios. We leave the design of more sophisticated KD framework specifically tailored for camera-radar fusion for future work.